OpenStudy (anonymous):

Solve on the interval [0, 2Pi): 1 + cos(theta) = (sqrt(3)+2)/2 A. Pi/6 , 5Pi/6 B. Pi/3 , 5Pi/3 C.Pi/6 , 11Pi/6 D. 7Pi/6, 11Pi/6 Please answer with explanation.

OpenStudy (anonymous):

it is the same as cos(theta)=sqrt(3)/2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

eh yeah i suppose ...why is it that ...what about the 1 and the 2? ...eh...\[(\sqrt{3} + 2) \div 2\]

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

cos(theta)=(sqrt(3)+2)/2 -1=sqrt(3)+2/2 - 2/2

OpenStudy (anonymous):

((3√+2)÷2) -1 not 2 -1

OpenStudy (anonymous):

\[\frac{ \sqrt{3}+2 }{ 2 }-1?\]

OpenStudy (anonymous):

thats what it would be

OpenStudy (nubeer):

\[\frac{ \sqrt{3} }{ 2 } +\frac{ 2 }{ 2 } -1\] \[\frac{ \sqrt{3} }{ 2 } +1 -1\] cos(theta) =√3 /2

OpenStudy (anonymous):

then what do i do? ..i understand that ty nu

OpenStudy (nubeer):

\[\theta = \cos^{-1} \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (nubeer):

as theta value would be positive means it lies in the 1st and 4th quadrant bcos in these quadrants value of cos is positive.

OpenStudy (anonymous):

then solve for \[\cos( \theta )=\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

wait ...can you reexplain how you moved cosine over?

OpenStudy (nubeer):

well its simple.. when u move cosine or any trignometric on other side of equal u have to inverse it.. do not mistake it for power ^-1 it is the inverse function.. u can find in calculator if u r allowed to use it.

OpenStudy (anonymous):

it's arccosin, not cosin

OpenStudy (anonymous):

arccosin is the inverse of cosin

OpenStudy (anonymous):

inverse function ?

OpenStudy (nubeer):

yes i think that's what is called.. they are also named as "arc" functions.

OpenStudy (anonymous):

sin^-1? thats 2nd sin button

OpenStudy (nubeer):

hmm yes maybe, it differs from calculator to calculator.. you have to find for cos^-1 on ur calculator and find the value can u tell me what u have got?

OpenStudy (anonymous):

be careful, there are some default setting in your calculator that it would not return all the results you want.

OpenStudy (anonymous):

it looks just like that ....and it depends on radians and degrees and stuff? if i put cos^-1 without anything it would be an error ..i need a value to put in it

OpenStudy (nubeer):

hmm by defualt it give in degree but u can change settings to get answer in radians.. and to find value do like this cos^-1 (√3/2) .. i think this will give u the value 30 degree.

OpenStudy (anonymous):

be careful of your interval. @Mpost1994

OpenStudy (anonymous):

yes and after that? eh..how do i get the answer?

OpenStudy (nubeer):

hmm just press equal sing on ur calculator u should get soem answer.

OpenStudy (anonymous):

no i meant for the question

OpenStudy (nubeer):

ohh sorry.. well u do u will get 30 or pi/6 in radians. as the value is positive . u have to see where cos function is positive .. it is in 1st Quadrant and 4th so , 1st quad = pi/6 4th quad = 2pi - pi/6

OpenStudy (anonymous):

11pi/6?

OpenStudy (nubeer):

yes.

OpenStudy (anonymous):

ok ty :3

OpenStudy (nubeer):

yw :)